OpenStudy (anonymous):

for f(x)=1/x^2-3 find for (2+h)

OpenStudy (misty1212):

hi!!

OpenStudy (misty1212):

\[f(\spadesuit)=\frac{1}{\spadesuit^2-3}\] replace \(\spadesuit\) by \(2+h\)

OpenStudy (anonymous):

@misty1212 i did I got 1/3h is that right?

OpenStudy (misty1212):

idk i did't do it, but i am sure it is wrong

OpenStudy (misty1212):

lets go slow

OpenStudy (misty1212):

first step is \[f(2+h)=\frac{1}{(2+h)^2-3}\]

OpenStudy (misty1212):

second step is to square the \(2+h\) i.e. do \((2+h)(2+h)=4+4h+h^2\) giving \[\frac{1}{4+4h+h^2-3}\]

OpenStudy (misty1212):

third step is the arithmetic, final answer is \[\frac{1}{4h+h^2+1}\]

OpenStudy (anonymous):

ohh ok thank you I forgot to square it, can you help me with one more problem?

OpenStudy (misty1212):

sure i will go to that post